# Grappa - A Machine Learned Molecular Mechanics Force Field

### v1

@@ A.2.2 Symmetric transformer
we apply $L_{\text{T}}$ transformer layers, each of which is given by scaled dot-product attention with $N_{\text{T-heads}}$ heads and a 2-layer MLP with hidden dimension $4d_{T}$, skip connections, dropout and layer normalization as in the original transformer architecture [7].
@@ A.2.3 Scaling of neural network outputs
To map the scores predicted from our model to the range of the respective MM parameter, we use scaled and shifted versions of the sigmoid function for mapping to a specified interval and of ELU for mapping to positive values. We choose these functions because they are differentiable and, in contrast to the exponential, only grow linearly for large inputs, which can help to stabilize training. To allow the model to predict scores that are of order unity, or approximately normally distributed, we choose the scaling and shifting parameters in such a way that a normally distributed input would lead to an output distribution whose mean and standard deviation are close to a given target mean $\mu$ and standard deviation $\sigma$. We choose this target standard deviation by calculating mean and standard deviation of traditional MM force field parameters on a given subset of the training data, except for equilibrium angles, where we restrict $\mu$ to $\pi/2$.
@@ A.3 Hyperparameters and training details
For all models discussed in this work, we use the hyperparameters listed in Table A3 and a dihedral periodicity of $n_{\text{periodicity}}=3$. All relative weights between energies, forces and MM parameters rely on units formed by kcal/mol, Å and radian. We train the models using the Adam [31] optimizer with $\epsilon=10^{-8}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$.
First, we train for 2 epochs on traditional MM parameters only, which can be interpreted as a form of pretraining that is efficient in bringing the model in a state where the gradients of the QM-part of the loss functions are informative. We start the training on energies and forces with $\lambda_{\text{trad}}=10^{-3}$, which we set to zero after 100 epochs, preventing the model from converging towards unphysical local minima of the QM-part of the loss function. When changing weights of terms of the loss function as described above, we re-initialize the optimizer with random moments and apply a warm restart [32], during which we linearly increase the learning rate in 500 training steps to decorrelate the optimizer state from the gradient updates of the previous loss function. Besides pretraining on traditional MM parameters directly, we have found penalizing large torsion and improper force constants $\xi^{(\text{dih})}$ to benefit generalizability across conformational space.
As validation loss, we use a linear combination of the energy and force root mean squared error (RMSE) on the sub-datasets (the rows in Table 1), averaged over all sub-datasets, with a weight of 1 for the energy RMSE and 3 for the force RMSE to prevent overfitting on a molecule type. We use this validation loss for early stopping and for learning rate scheduling. For the model trained on the Espaloma dataset (Table 1), the validation loss did not improve after 1000 epochs, which corresponds to one day of training on a single NVIDIA RTX 6000 Ada GPU.
@@ A.4 Improper dihedrals
which can be seen from the formula for the dihedral angle in [33]. This symmetry allows us to identify pairs of terms that are transformed into each other under the permutation $ijkl\rightarrow ljki$ and use a force constant that is invariant under this permutation,
@@ A.5.1 Espaloma dataset
We reproduce the dataset used to train and evaluate Espaloma 0.3 [10] by downloading the published, preprocessed data [^2] https://github.com/hits-mbm-dev/grappa/blob/master/dataset_creation/get_espaloma_split/load_esp_ds.py. The preprocessing entails filtering out high-energy conformations as detailed in [10]. As in Espaloma, our train-val-test partitioning procedure relies on the unique isomeric SMILES strings of the molecules, which are included in the dataset. We reproduce [^3] https://github.com/hits-mbm-dev/grappa/blob/master/dataset_creation/get_espaloma_split/load_esp_ds.py the partitioning from Espaloma 0.3444https://github.com/choderalab/refit-espaloma/blob/main/openff-default/02-train/joint-improper-charge/charge-weight-1.0/train.py, commit 3ccc44d with the same random seed.
@@ A.5.2 Dataset creation
Scripts for the creation of the datasets can be found at https://github.com/LeifSeute/grappa-data-creation. We use the Amber ff99SB-ILDN [17] force field for sampling the states and Psi4 [34] for the single point QM calculations. For sampling radical peptide states, we use a preliminary Grappa model trained on optimization trajectories and torsion scans of radical peptides. The radical peptides are what we call hydrogen-atom-transfer (HAT) type radicals, that is, they are formed by removing a single hydrogen from a peptide. For the nonbonded contribution, we use the same Lennard-Jones parameters as for the original peptide and add the partial charge of the hydrogen in the original peptide to the heavy atom it was attached to.
@@ A.6 MD simulations
All folding and stability simulations are performed with GROMACS version 2023 [8] using the Amber ff99SB-ILDN force field [17] or Grappa-1.1 with Amber ff99SB-ILDN nonbonded parameters and the TIP3P water model. 2 fs time steps with LINCS constraints [35] on H-bonds were employed. Simulations were run at 300K and 1 bar, maintained by the v-rescale thermostat [36] and Parrinello-Rahman pressure coupling [37]. We apply a Coulomb and Lennard-Jones cutoff of 1.0ṅm. After system preparation, energy minimization, NVT and NPT equilibration simulations are conducted. The code for reproducing the simulations can be found at
https://github.com/LeifSeute/validate-grappa.
A.6.1 Folding simulations
For the folding simulations, we follow the simulation system setup of Lindorff-Larsen et al. [25] From a completely extended solvated structure, we minimize and equilibrate, followed by a short MD simulation, in which the end-to-end distance of the protein usually reduces. We pick a state with low end-to-end distance and use it as starting point for the actual folding simulation in which we can choose the simulation box smaller, reducing the computational cost. For the folding simulations of a mutant of chignolin [38] (PDB ID: 5AWL) with the sequence YYDPETGTWY, we choose a box size of 40 Å, and neutralize the system with two sodium ions.
A.6.2 Stability simulations
The simulations of ubiquitin (PDB ID: 1UBQ) are evaluated by calculating a moving average over 0.05 ns of the C-alpha RMSD to the initial structure (Figure 4b). We also calculate the statistics of the C-alpha RMSD as function of time difference between two states. To this end, we sample 1,000 states from the trajectory for each time difference and calculate the mean C-alpha RMSD between the states.
Figure A9: Test RMSE of Grappa (blue), trained on a fraction of the Espaloma train dataset, compared with the RMSE of Espaloma (green) trained on the full dataset and the RMSE of established force fields.
@@ A.6.2 Stability simulations
Table A4: Accuracy of Grappa, Espaloma 0.3 [10] and established MM force fields on test molecules from the Espaloma dataset. We report the RMSE of molwise-centered energies in kcal/mol and the componentwise RMSE of forces in kcal/mol/Å. Gaff-2.11 [21] is a general-purpose force field, ff14SB [22] is an established protein force field and RNA.OL3 [23] is specialized to RNA. As in Espaloma, we bootstrap the set of test molecules 1000 times and report mean and standard deviation of the RMSEs.
Dataset Test Mols Confs Grappa Grappa Espaloma ff14SB, Mean
AM1-BCC ff99SB RNA.OL3 Predictor
BOLTZMANN SAMPLED
SPICE-Pubchem 1411 60853 Energy 2.3 2.3 18.6
Force 6.1 6.8 23.5
SPICE-DES-Monomers 39 2032 Energy 1.3 1.4 8.3
Force 5.3 5.9 21.3
SPICE-Dipeptide 67 2592 Energy 2.3 2.2 3.1 4.6 4.3
Force 5.5 5.9 7.8 12.1 10.7
RNA-Diverse 6 357 Energy 3.4 4.2 6.0 5.8
Force 3.7 4.4 19.4 17.0
RNA-Trinucleotide 64 35811 Energy 3.6 3.8 6.1 5.8
Force 3.6 4.3 19.7 17.8
Radical-Dipeptide 28 275 Energy 2.8 7.4
Force 6.5 23.6
HYP-DOP-Dipeptide 8 363 Energy 2.3 6.8
Force 5.6 24.9
Uncapped-Dipeptide 12 600 Energy 2.1 3.7 6.5
Force 5.9 11.7 25.3
@@
Gen2-Torsion 131 21890 Energy 1.8 1.6 4.8
Force 4.0 4.7 5.5
Protein-Torsion 9 6624 Energy 2.2 2.2 1.9 2.4
Force 3.9 5.2 3.5 9.6
OPTIMIZATION
Gen2-Opt 154 40055 Energy 1.8 1.7 3.9
Force 3.8 4.5 5.1
Pepconf-Opt 55 14884 Energy 3.2 2.7 2.8 4.1 3.5
Force 3.7 4.5 4.0 10.2 9.6
Table A5: Accuracy of Grappa with nonbonded parameters from the AM1-BCC scheme, Grappa with nonbonded parameters from Amber ff99SB-ILDN, Espaloma 0.3 [10] and established MM force fields on test molecules of the Grappa-1.1 dataset. We report the RMSE of molwise-centered energies in kcal/mol and the componentwise RMSE of forces in kcal/mol/Å.

### v2

@@ A.2.2 Symmetric transformer
we apply $L_{\text{T}}$ transformer layers, each of which is given by scaled dot-product attention with $N_{\text{T-heads}}$ heads and a 2-layer MLP with hidden dimension $4d_{T}$, skip connections, dropout and layer normalization as in the original transformer architecture (Vaswani et al., 2023).
@@ A.2.3 Scaling of neural network outputs
To map the scores predicted from the model to the range of the respective MM parameter, we use scaled and shifted versions of the sigmoid function for mapping to a specified interval and of ELU for mapping to positive values. We choose these functions because they are differentiable and, in contrast to the exponential, only grow linearly for large inputs, which can help to stabilize training. To allow the model to predict scores that are of order unity, or approximately normally distributed, we choose the scaling and shifting parameters in such a way that a normally distributed input would lead to an output distribution whose mean and standard deviation are close to a given target mean $\mu$ and standard deviation $\sigma$. We choose this target standard deviation by calculating mean and standard deviation of traditional MM force field parameters on a given subset of the training data, except for equilibrium angles, where we restrict $\mu$ to $\pi/2$.
@@ A.3 Hyperparameters and training details
For all models discussed in this work, we use the hyperparameters listed in Table A5 and a dihedral periodicity of $n_{\text{periodicity}}=3$. All relative weights between energies, forces and MM parameters rely on units formed by kcal/mol, Å and radian. We train the models using the Adam (Kingma & Ba, 2017) optimizer with $\epsilon=10^{-8}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$.
First, we train for 2 epochs on traditional MM parameters only, which can be interpreted as a form of pretraining that is efficient in bringing the model in a state where the gradients of the QM-part of the loss functions are informative. We start the training on energies and forces with $\lambda_{\text{trad}}=10^{-3}$, which we set to zero after 100 epochs, preventing the model from converging towards unphysical local minima of the QM-part of the loss function. When changing weights of terms of the loss function as described above, we re-initialize the optimizer with random moments and apply a warm restart (Loshchilov & Hutter, 2017), during which we linearly increase the learning rate in 500 training steps to decorrelate the optimizer state from the gradient updates of the previous loss function. Besides pretraining on traditional MM parameters directly, we have found penalizing large torsion and improper force constants $\xi^{(\text{dih})}$ to benefit generalizability across conformational space.
As validation loss, we use a linear combination of the energy and force root mean squared error (RMSE) on the sub-datasets (the rows in Table 1), averaged over all sub-datasets, with a weight of 1 for the energy RMSE and 3 for the force RMSE to prevent overfitting on a molecule type. We use this validation loss for early stopping and for learning rate scheduling. For the model trained on the Espaloma dataset (Table 1), the validation loss did not improve after 1000 epochs, which corresponds to about 25 hours of training on an A100 GPU.
@@ A.4 Improper dihedrals
which can be seen from the formula for the dihedral angle in (Blondel & Karplus, 1996). This symmetry allows us to identify pairs of terms that are transformed into each other under the permutation $ijkl\rightarrow ljki$ and use a force constant that is invariant under this permutation,
@@ A.5.1 Espaloma dataset
We reproduce the dataset used to train and evaluate Espaloma 0.3 (Takaba et al., 2023) by downloading the published, preprocessed data666 https://github.com/graeter-group/grappa/blob/master/dataset_creation/get_espaloma_split/load_esp_ds.py . The preprocessing entails filtering out high-energy conformations as detailed in (Takaba et al., 2023). As in Espaloma, our train-val-test partitioning procedure relies on the unique isomeric SMILES strings of the molecules, which are included in the dataset. We reproduce777 https://github.com/graeter-group/grappa/blob/master/dataset_creation/get_espaloma_split/load_esp_ds.py the partitioning from Espaloma 0.3888 https://github.com/choderalab/refit-espaloma/blob/main/openff-default/02-train/joint-improper-charge/charge-weight-1.0/train.py, commit 3ccc44d with the same random seed.
@@ A.5.2 Dataset creation
Scripts for the creation of the datasets are publicly available. 999 https://github.com/LeifSeute/grappa-data-creation We use the Amber ff99SB-ILDN (Lindorff-Larsen et al., 2010) force field for sampling the states and Psi4 (Smith et al., 2020) for the single point QM calculations. For sampling radical peptide states, we use a preliminary Grappa model trained on optimization trajectories and torsion scans of radical peptides. The radical peptides are what we call hydrogen-atom-transfer (HAT) type radicals, that is, they are formed by removing a single hydrogen from a peptide. For the nonbonded contribution, we use the same Lennard-Jones parameters as for the original peptide and add the partial charge of the hydrogen in the original peptide to the heavy atom it was attached to.
@@ A.6 MD simulations
All molecular dynamics simulations are performed with GROMACS version 2023 (Abraham et al., 2015) using the Amber ff99SB-ILDN force field (Lindorff-Larsen et al., 2010) or Grappa-1.3 with Amber ff99SB-ILDN nonbonded parameters and the TIP3P water model. 2 fs time steps with LINCS constraints (Hess, ) on H-bonds were employed. Simulations were run at 300K and 1 bar, maintained by the v-rescale thermostat (Bussi et al., ) and Parrinello-Rahman pressure coupling (Parrinello & Rahman, ). We apply a Coulomb and Lennard-Jones cutoff of 1.0ṅm. After system preparation, energy minimization, NVT and NPT equilibration simulations are conducted. The code for reproducing the simulations is publicly available. 101010 https://github.com/LeifSeute/validate-grappa.
A.6.1 Ubiquitin simulations
The simulations of ubiquitin (PDB ID: 1UBQ) are evaluated by calculating a moving average over 0.05 ns of the C-alpha RMSD to the initial structure (Figure 5b). We also calculate the statistics of the C-alpha RMSD as function of time difference between two states. To this end, we sample 1,000 states from the trajectory for each time difference and calculate the mean C-alpha RMSD between the states. For the simulation of STMV, we use the same structure files as (Musaelian et al., 2022). We use partial charges and Lennard-Jones parameters of Amber ff99SB-ILDN (Lindorff-Larsen et al., 2010) for the proteins and of RNA.OL3 (Zgarbová et al., 2011) for RNA.
A.6.2 STMV virus
For the simulation of STMV referred to in section 4.2, we use nonbonded parameters from Amber ff99SB and RNA.OL3 repsectively.
Fig. A9 (b) shows the RMSD of protein C-alpha and RNA carbon atoms during 30 ns of MD simulation with Grappa parameters for protein and RNA, which indicates that Grappa keeps the virus stable. For large scale conformational changes or artificial viral capsid disruption, higher RMSD values would be expected.
Figure A9: (a) The virus STMV, containing proteins (blue) and RNA (red). (b) RMSD of protein C-alpha and RNA carbon atoms of the virus in solution during a 30 ns MD simulation with Grappa.
A.7 Learning curve
We provide a learning curve for Grappa, trained on subsets of the Espaloma dataset, in Fig. A10.
A.8 Grappa’s MM parameters
In Fig. A11, the MM parameters predicted by Grappa and ff99-SBILDN for the protein ubiquitin are visualized.
A.9 Force accuracy
In analogy to Fig. 4, Grappa’s accuracy for predicted force components is visualized in Fig. A12.
A.10 Extensive tables
We provide tables for with errors for the full Grappa-1.3 and Espaloma datasets in Tab. A6 and A7.
Figure A10: Test RMSE of Grappa (blue), trained on a fraction of the Espaloma-0.3 train dataset, compared with the RMSE of Espaloma-0.3 (green) trained on the full dataset and the RMSE of established force fields.
Figure A11: MM parameters for the protein ubiquitin, predicted by Grappa-1.3 and ff99SB-ILDN. The comparison suggests that Grappa predicts more continuous parameter sets than the tabulated force field ff99SB-ILDN.
Figure A12: Comparison of force predictions of (a) Grappa-1.3 and the established force fields (b) Gaff-2.11, ff99SB-ILDN and RNA.OL3 for test molecules from Espaloma’s SPICE-Pubchem, SPICE-Dipeptide and RNA-Trinucleotide datasets. For SPICE-PubChem, five molecules were filtered out because they had unphysically high gradient magnitudes of more than 400 kcal/mol/Å.
@@ A.10 Extensive tables
Table A6: Accuracy of Grappa, Espaloma 0.3 (Takaba et al., 2023) and established MM force fields on test molecules from the Espaloma dataset. We report the RMSE of molwise-centered energies in kcal/mol and the componentwise RMSE of forces in kcal/mol/Å. Gaff-2.11 (He et al., 2020) is a general-purpose force field, ff14SB (Maier et al., 2015) is an established protein force field and RNA.OL3 (Zgarbová et al., 2011) is specialized to RNA. As in Espaloma, we bootstrap the set of test molecules 1000 times and report mean and standard deviation of the RMSEs.
Dataset Test Mols Confs Grappa Espaloma Gaff-2.11 ff99SB-ILDN Mean
Predictor
SPICE-PubChem 1411 60853 Energy 2.3 $\pm$ 0.1 2.3 $\pm$ 0.1 4.6 $\pm$ 0.1 18.8
Force 6.1 $\pm$ 0.3 6.8 $\pm$ 0.1 14.6 $\pm$ 0.3 41.3
SPICE-DES-Monomers 39 2032 Energy 1.3 $\pm$ 0.1 1.4 $\pm$ 0.3 2.5 $\pm$ 0.2 8.6
Force 5.3 $\pm$ 0.2 5.9 $\pm$ 0.5 11.1 $\pm$ 0.6 37.4
SPICE-Dipeptide 67 2592 Energy 2.4 $\pm$ 0.1 3.1 $\pm$ 0.1 4.5 $\pm$ 0.1 19.1
Force 5.4 $\pm$ 0.1 7.8 $\pm$ 0.2 12.9 $\pm$ 0.3 38.4
RNA-Diverse 6 357 Energy 3.2 $\pm$ 0.2 4.2 $\pm$ 0.3 6.5 $\pm$ 0.1 6.1
Force 3.7 $\pm$ 0.0 4.4 $\pm$ 0.1 16.8 $\pm$ 0.1 30.2
RNA-Trinucleotide 64 23811 Energy 3.5 $\pm$ 0.0 3.8 $\pm$ 0.2 6.0 $\pm$ 0.1 6.1
Force 3.6 $\pm$ 0.0 4.3 $\pm$ 0.1 17.0 $\pm$ 0.0 31.2
Dipeptides-300K 72 3600 Energy 2.6 $\pm$ 0.1 4.5 $\pm$ 0.1 4.1 $\pm$ 0.1 7.8
Force 5.9 $\pm$ 0.1 10.9 $\pm$ 0.3 11.8 $\pm$ 0.1 43.7
Dipeptides-1000K 72 2160 Energy 5.4 $\pm$ 0.1 8.6 $\pm$ 0.2 8.5 $\pm$ 0.2 18.9
Force 11.6 $\pm$ 0.1 16.1 $\pm$ 0.2 17.8 $\pm$ 0.1 74.3
Non-Capped-Peptides 10 500 Energy 2.2 $\pm$ 0.2 4.2 $\pm$ 0.4 4.0 $\pm$ 0.3 7.2
Force 6.1 $\pm$ 0.2 12.1 $\pm$ 0.7 12.6 $\pm$ 0.5 45.8
Radical-Dipetides 28 272 Energy 3.3 $\pm$ 0.3 8.7
Force 6.8 $\pm$ 0.2 41.3
OPTIMIZATION
Gen2-Opt 154 29055 Energy 1.7 $\pm$ 0.1 1.7 $\pm$ 0.5 2.8 $\pm$ 0.2 4.2
Force 4.0 $\pm$ 0.2 4.5 $\pm$ 0.8 9.8 $\pm$ 0.3 8.7
Pepconf-Opt 55 9084 Energy 2.8 $\pm$ 0.2 2.8 $\pm$ 0.3 4.7 $\pm$ 0.3 6.5
Force 3.7 $\pm$ 0.2 4.0 $\pm$ 0.4 10.4 $\pm$ 0.3 9.4
@@
Gen2-Torsion 131 19290 Energy 1.7 $\pm$ 0.1 1.6 $\pm$ 0.3 2.6 $\pm$ 0.1 4.7
Force 4.2 $\pm$ 0.2 4.7 $\pm$ 0.6 9.5 $\pm$ 0.4 8.9
Protein-Torsion 9 6024 Energy 1.9 $\pm$ 0.2 1.9 $\pm$ 0.2 2.9 $\pm$ 0.2 3.5
Force 4.2 $\pm$ 0.3 3.5 $\pm$ 0.3 9.8 $\pm$ 0.4 9.0
Table A7: Accuracy of Grappa-1.3 with nonbonded parameters from the AM1-BCC scheme, Grappa with nonbonded parameters from Amber ff99SB-ILDN, Espaloma 0.3 (Takaba et al., 2023) and established MM force fields on test molecules of the Grappa-1.3 dataset. We report the RMSE of molwise-centered energies in kcal/mol and the componentwise RMSE of forces in kcal/mol/Å.